OpenStudy (anonymous):

perform the indicated function evaluations for: f(x)=10x-3 f(12-x)....?

jimthompson5910 (jim_thompson5910):

Replace 'x' with '12-x' to get f(x)=10x-3 f(12-x)=10(12-x)-3 Now simplify and tell me what you get.

OpenStudy (anonymous):

120x-3 .........?

jimthompson5910 (jim_thompson5910):

10(12-x) - 3 10*12-10*x - 3 120 - 10x - 3 117x - 10x -10x + 117 So f(12-x)= -10x+117

jimthompson5910 (jim_thompson5910):

So somehow you mixed up the terms

OpenStudy (anonymous):

soo I probably messed up on f(t^2+2) ....

jimthompson5910 (jim_thompson5910):

perhaps, but idk since I can't see what you got

OpenStudy (anonymous):

I got 10t^2+20-3 for that one

jimthompson5910 (jim_thompson5910):

f(x)=10x-3 f(t^2+t)=10(t^2+t)-3 f(t^2+t)=10t^2+10t-3

jimthompson5910 (jim_thompson5910):

and you can't go further because there are no like terms to combine

OpenStudy (anonymous):

my problem with these types of equations is that I always end up multiplying whats outside of the parenthesis with whats inside

OpenStudy (anonymous):

thats how I got 20 instead of 10

jimthompson5910 (jim_thompson5910):

well you're sorta on the right track because you do multiply that 10 sitting out front the parenthesis with everything inside

OpenStudy (anonymous):

so why wouldn't I multiply it with the 2 thats sitting inside the parenthesis?

jimthompson5910 (jim_thompson5910):

because you only multiply the coefficients, not the exponents

jimthompson5910 (jim_thompson5910):

since 10 times t^2 = 10t^2 and NOT 20t^2, 20t or 20

OpenStudy (anonymous):

so just ignore the 2 right

jimthompson5910 (jim_thompson5910):

pretty much

OpenStudy (anonymous):

ok thankyou so much!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

ok so would f(x-h) be 10x+10h-3...?

jimthompson5910 (jim_thompson5910):

f(x)=10x-3 f(x-h)=10(x-h)-3 Got this so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh wait I made a mistake its supposed to be a plus sign not a minus

jimthompson5910 (jim_thompson5910):

Now distribute f(x)=10x-3 f(x-h)=10(x-h)-3 f(x-h)=10x-10h-3

OpenStudy (anonymous):

yeah thats what I got except its a plus sign between 10x and 10h

jimthompson5910 (jim_thompson5910):

i see, so you lost a negative sign somewhere

OpenStudy (anonymous):

do you mind helping me with another one real fast..?

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

ok its sorta the same kind of problem, h(y)=4y^2-7y+1 h(6z)...?

jimthompson5910 (jim_thompson5910):

Start with h(y)=4y^2-7y+1 and replace EVERY 'y' with '6z' to get h(6z)=4(6z)^2-7(6z)+1 Then simplify

OpenStudy (anonymous):

12z^2-42z+1

jimthompson5910 (jim_thompson5910):

(6z)^2 = 36z^2 then 4*36z^2 = 144z^2 So it should be 144z^2-42z+1

OpenStudy (anonymous):

I thought after I did that I would then square 12 which would be 144

jimthompson5910 (jim_thompson5910):

how are you getting 12?

jimthompson5910 (jim_thompson5910):

or 12^2 ?

jimthompson5910 (jim_thompson5910):

sure 12^2 = 144 and 4*36 = 144, but I'm not seeing how you got 12

OpenStudy (anonymous):

yeap you're right I was multiplying wrong

jimthompson5910 (jim_thompson5910):

i gotcha, i'm glad you caught your mistake

OpenStudy (anonymous):

haha yeah. ok I'm gonna work one out on my own and you tell me if I get it right?

jimthompson5910 (jim_thompson5910):

alright sounds great

OpenStudy (anonymous):

cool! ok this one looks hard ...but I'll give it a try... its for the same problem .... h(1-3y) 4(1-3y)^2-7(1-3y)+1= 4-12y^2-7-21+1

OpenStudy (anonymous):

and Idk if I should keep simplifying...?

jimthompson5910 (jim_thompson5910):

it might help to just focus on (1-3y)^2 (1-3y)^2 (1-3y)(1-3y) 1(1-3y)-3y(1-3y) So simplify 1(1-3y)-3y(1-3y) and tell me what you get

OpenStudy (anonymous):

3y-3y-9y^2

OpenStudy (anonymous):

Ik thats not right...BUtik you have to multiply it with whats int the parenthesis..

jimthompson5910 (jim_thompson5910):

1(1-3y)-3y(1-3y) 1-3y-3y-9y^2 1 - 6y - 9y^2 Somehow you lost a 1

OpenStudy (anonymous):

YEAH I did. I make the littlest mistakes -__- Ok where did the 6 some from?

jimthompson5910 (jim_thompson5910):

-3y - 3y = -6y

jimthompson5910 (jim_thompson5910):

since -3-3 = -6

OpenStudy (anonymous):

ok I get it now

jimthompson5910 (jim_thompson5910):

alright great

jimthompson5910 (jim_thompson5910):

so 4(1-3y)^2-7(1-3y)+1 becomes 4(1 - 6y - 9y^2)-7(1-3y)+1

OpenStudy (anonymous):

thankyou!

jimthompson5910 (jim_thompson5910):

you're welcome, tell me what you get when you simplify that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry for the wait. ok so I got 4-24y-81y^2-7-21-1...?

jimthompson5910 (jim_thompson5910):

you mean -36y^2 ? and -21y?

jimthompson5910 (jim_thompson5910):

oh and that -1 at the end should be +1

jimthompson5910 (jim_thompson5910):

oh and sry, -7 times -3y is +21y

jimthompson5910 (jim_thompson5910):

let me write it all out 4(1 - 6y - 9y^2)-7(1-3y)+1 4 - 24y - 36y^2 - 7 + 21y + 1

jimthompson5910 (jim_thompson5910):

now combine like terms

OpenStudy (anonymous):

4-36y^2-7+45y+1

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

-24+21 = -3

jimthompson5910 (jim_thompson5910):

so -24y + 21y = -3y

jimthompson5910 (jim_thompson5910):

4-7+1 = -3+1 = -2

jimthompson5910 (jim_thompson5910):

4 - 24y - 36y^2 - 7 + 21y + 1 -36y^2 - 3y - 2

OpenStudy (anonymous):

Does it matter which one I subtracted or added first?

jimthompson5910 (jim_thompson5910):

with subtraction order matters, so yes

jimthompson5910 (jim_thompson5910):

-24+21 is the same as 21 - 24

OpenStudy (anonymous):

ok...I just got lost.....

jimthompson5910 (jim_thompson5910):

where

jimthompson5910 (jim_thompson5910):

point which step

jimthompson5910 (jim_thompson5910):

point out*

OpenStudy (anonymous):

at this point...4-7+1 = -3+1 = -2

jimthompson5910 (jim_thompson5910):

4 - 7 is -3

jimthompson5910 (jim_thompson5910):

so 4-7+1 is the same as -3+1

jimthompson5910 (jim_thompson5910):

then -3+1 becomes -2

jimthompson5910 (jim_thompson5910):

So in the end, 4-7+1 = -2

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

Here's the full and complete problem 4(1-3y)^2-7(1-3y)+1 4[ (1-3y)(1-3y) ]-7(1-3y)+1 4(1 - 6y + 9y^2)-7(1-3y)+1 4 - 24y + 36y^2 - 7 + 21y + 1 36y^2 - 3y - 2

jimthompson5910 (jim_thompson5910):

So 4(1-3y)^2-7(1-3y)+1 simplifies to 36y^2 - 3y - 2

OpenStudy (anonymous):

how did you get 36y^2-3y-2 as your final answer..... I see how you got 36....but where did 3 and 2 come from...

jimthompson5910 (jim_thompson5910):

Hopefully this is a bit clearer 4(1-3y)^2-7(1-3y)+1 4[ (1-3y)(1-3y) ]-7(1-3y)+1 4(1 - 6y + 9y^2)-7(1-3y)+1 4 - 24y + 36y^2 - 7 + 21y + 1 36y^2 (- 24y+21y) + (4- 7+ 1 ) 36y^2 - 3y - 2

jimthompson5910 (jim_thompson5910):

If not, let me know

OpenStudy (anonymous):

got it! makes sense

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

thankyou so much ! you're a great teacher.

jimthompson5910 (jim_thompson5910):

why thank you, I'm glad I could help you out

OpenStudy (anonymous):

is it ok if you help me with a couple more problems. Im trying to get the hang of this.

jimthompson5910 (jim_thompson5910):

sure go for it

OpenStudy (anonymous):

this one is h(y+k) I got 4(y+k)^2-7(y-K)+1 =4(y+k)^2-7(y+7yk+k)+1 =4y^2+4k^2-7y+49yk+7k+1

jimthompson5910 (jim_thompson5910):

h(y+k) ?? not f(y+k) ??

jimthompson5910 (jim_thompson5910):

wait, we're talking about the same function 4x^2 - 7x + 1 right??

OpenStudy (anonymous):

nooo its h

jimthompson5910 (jim_thompson5910):

oh ok, but it's still 4x^2 - 7x + 1 right?

OpenStudy (anonymous):

yesss

jimthompson5910 (jim_thompson5910):

ok thanks